OpenStudy (anonymous):

The height of a parallelogram is 2 units more than the base. The measure of the base of the parallelogram is 10 units. What is the area, in square units, of the parallelogram? 200 60 100 120

OpenStudy (anonymous):

\[Base \times Base +2\]

OpenStudy (anonymous):

The area of a parallelogram is A=BH in this case your height is two more than your base.

OpenStudy (lgbasallote):

@angelabink you wrote that notation wrong heh..according to PEMDAS you should write it as \[\text{base} \times (\text{base} + 2)\] ;)

OpenStudy (anonymous):

i keep getting 240 and ik that that is wrong

OpenStudy (cwrw238):

base = 10, height = 10 + 2 area = height * base

OpenStudy (anonymous):

ok thanks

OpenStudy (cwrw238):

so what do you think answer is?

OpenStudy (anonymous):

120

OpenStudy (cwrw238):

right